OpenStudy (anonymous):

a man got a 4000 dollar said which equated to an 11% raise. What was his salary before the raise?

OpenStudy (anonymous):

i set it up .11(x) = 4000+x

OpenStudy (anonymous):

i got roughly 36,000

OpenStudy (anonymous):

it should say got a 4000 dollar raise which equated to 11% increase in his salary*****

OpenStudy (anonymous):

so you are asking $4000 is 11% of what number

OpenStudy (anonymous):

translate into math \[4000=.11\times x\]

OpenStudy (anonymous):

"is" becomes = "of" becomes x and "what number" is our variable.

OpenStudy (anonymous):

solve in one step via \[x=\frac{4000}{.11}\]

OpenStudy (anonymous):

oh and of course i should add that 11% becomes .11 because we compute with numbers, not percents.

OpenStudy (anonymous):

i get an original salary of $36,363.63

OpenStudy (anonymous):

cool, thats what I got i actually had 4000-x = 1.11*x

OpenStudy (anonymous):

so it equated to the same but by dumb luck :) thanks!

OpenStudy (anonymous):

i mean 4000+x wow im struggling

OpenStudy (anonymous):

well actually your equation was exactly the same i think. just more work to solve and rather convoluted for this basic problem

OpenStudy (anonymous):

hahah yes I agree

OpenStudy (anonymous):

subtract x, get \[.11x=4000\] etc